### v1

# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

Abstract.
We prove Vojta’s abc conjecture for projective space ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming that the entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$ are highly ramified over the coordinate hyperplanes. This extends the results of Guo Ji and the second-named author for the case $n=2$ (see [9]). We also explore the corresponding results for projective toric varieties. Consequently, we establish a version of Campana’s orbifold conjecture for finite coverings of projective toric varieties.
2020 Mathematics Subject Classification: Primary 30D35; Secondary 32Q45 and 32H30
The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604. The second-named author was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
1. Introduction
Let $X$ be an algebraic variety, and $D$ an effective divisor (possibly empty). The set $X\backslash D$ is Brody hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is constant. It is Brody quasi-hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is degenerate, meaning its image is contained a proper Zariski-closed subset $Z$ of $X$. We call such $Z$ the exceptional set. If $Z$ is independent of $f$, we call $X\backslash D$ Brody strong quasi-hyperbolic. This leads us to the Griffiths-Lang’s conjecture.
Griffiths-Lang’s conjecture. Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ the canonical divisor on $X$. If $K_{X}+D$ is big, then $X\backslash D$ is Brody quasi-hyperbolic.
A more precise quantitative version, in terms of the Nevanlinna’s functions, can be formulated as follows.
Griffiths’ conjecture. For a holomorphic map $f:\mathbb{C}\to X$ with Zariski dense image, the inequality
$T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}(D,r)+\epsilon T_{A,f}(r)$
holds for any $\epsilon>0$, where notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
Here, for each positive integer $n$, $N_{f}^{(n)}(D,r)$ is the $n$-truncated counting function with respect to $D$ given by
$N_{f}^{(n)}(D,r)=\sum_{0<|z|<r}\mathop{\mathrm{min}}\{{\rm ord}_{z}f^{*}D,n\}\log\frac{r}{|z|}+\mathop{\mathrm{min}}\{{\rm ord}_{0}f^{*}D,n\}\log r,$
$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [25, Section 12]).
Paul Vojta formulated the following general abc conjecture (see [25, Conjecture 15.2] and [25, Conjecture 23.4]), which extends the above Griffiths’ conjecture. It is also viewed as the quantitive version of Campana’s orbifold conjecture which will be stated later in the introduction.
Vojta’s abc conejcture. For any $\epsilon>0$, there exists a proper Zariski-closed subset $Z$ of $X$ such that for any entire curve $f:\mathbb{C}\to X$ whose image is not contained in $Z$, the inequality
$T_{K_{X}+D,f}(r)\leq_{\rm exc}N_{f}^{(1)}(D,r)+\epsilon T_{A,f}(r)$
holds, where $A$ is an ample divisor on $X$.
Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set is independent of the map $f$, and (ii) the counting function is truncated at level one. Proving these two features is quite challenging, even if the aforementioned version of Griffiths’ conjecture could be established. In this paper, we first prove Vojta’s abc conjecture for entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming the map is highly ramified over the coordinate hyperplanes. The precise statement of our main result is as follows. Additionally, we highlight that the exceptional set obtained in the theorem can be determined explicitly-this is a challenging issue in the study of Nevanlinna theory and complex hyperbolicity.
Main Theorem.
Let $G$ be a non-constant homogeneous polynomial in $\mathbb{C}[x_{0},\dots,x_{n}]$ with no monomial factors and no repeated factors. Assume that $[G=0]$ is in general position with the coordinate hyperplanes $H_{i}=[x_{i}=0]$ in $\mathbb{P}^{n}(\mathbb{C})$, $0\leq i\leq n$. Then, for any $\epsilon>0$, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $Z$ of $\mathbb{P}^{n}({\mathbb{C}})$ that can be explicitly determined, such that the following two inequalities hold:
  • (i)
    $N_{G(\mathbf{f})}(0,r)-N^{(1)}_{G(\mathbf{f})}(0,r)\leq_{\operatorname{exc}}\epsilon T_{f}(r)$, and
  • (ii)
    $(\deg G-\epsilon)T_{f}(r)\leq_{\operatorname{exc}}N^{(1)}_{G(\mathbf{f})}(0,r)$,
for any non-constant holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ whose image is not contained in $Z$, and where ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$ with $f(z_{0})\in H_{i}$. Here, the evaluation of $G$ is taken at $\mathbf{f}=(f_{0},\ldots,f_{n})$ in a reduced form of $f$, meaning that $f_{0},\ldots,f_{n}$ are entire functions without common zeros.
Note that the case when $n=2$ is due to Ji Guo and the second named author ([9, Theorem 1.4]). The new idea in this paper is motivated by the recent paper [2] by considering the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case. To achieve this, we need to extend the results in [9] and [7] to the parabolic setting.
We state some consequences of our Main Theorem, starting with the introduction of Campana’s orbifolds. Instead of considering the complement case in Brody’s hyperbolicity, we can also examine an entire curve $f:\mathbb{C}\to X$ that is ramified over $D$. According to Campana [1], an orbifold divisor is defined as $\Delta:=\sum_{Y\subset X}(1-m_{\Delta}^{-1}(Y))\cdot Y$, where $Y$ ranges over all irreducible divisors of $X$, and $m_{\Delta}(Y)\in[1,\infty)\cap{\mathbb{Q}}$, with all but finitely many values equal to one. An orbifold entire curve $f:{\mathbb{C}}\rightarrow(X,\Delta)$, with $\Delta=\sum_{i}(1-m_{i}^{-1})D_{i}$, is a homomorphic map such that $f({\mathbb{C}})\not\subset\mbox{Supp}(\Delta)$ and ${\rm mult}_{z_{0}}(f^{*}D_{i})\geq m_{i}$ for all $i$ and for every $z_{0}\in{\mathbb{C}}$ with $f(z_{0})\in D_{i}$.
Campana’s orbifold conjecture. If $(X,\Delta$) is an orbifold pair such that $K_{X}+\Delta$ is big, then there exists a proper closed subvariety $Z\subset X$ containing all the non-constant orbifold entire curves $f:{\mathbb{C}}\rightarrow(X,\Delta)$.
Remarks.
  1. (i)
    The Campana’s conjecture is a consequence of the above Vojta’s strong abc conecutre.
  2. (ii)
    Let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$. Recall that the pair $(X,D)$ is said to be of log-general type if the divisor $K_{X}+D$ is big. It’s clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e. $K_{X}+\Delta$ is big. On the other hand, the condition that $(X,D)$ is of log-general type implies that $(X,\Delta)$ is of general type if $(X,\Delta)$ has a sufficiently large multiplicity along each $D_{i}$, $1\leq i\leq q$. ((((See [12, Corollary 2.2.24].))))
The Main Theorem leads to the following two theorems concerning Campana’s orbifold conjecture; however, we will not include the proofs, as they can be adapted from those in [9] for the case $n=2$. We begin with the result concerning $n+2$ components.
Theorem 1.
Let $\Delta_{0}$ be an orbifold divisor of ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position. Let $m_{i}\in(1,\infty)\cap{\mathbb{Q}}$, $0\leq i\leq n$, and $\Delta=\Delta_{0}+(1-{1\over m_{0}})H_{0}+(1-{1\over m_{1}})H_{1}+\cdots+(1-{1\over m_{n}})H_{n}$. Assume that $\deg\Delta>n+1$. Then there exist a proper Zariski closed subset $W$ of ${\mathbb{P}}^{n}({\mathbb{C}})$ and an effectively computable positive integer $\ell$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{0},m_{1},\dots,m_{n}\}\geq\ell$ must be contained in $W$.
The next case involves $n+1$ components, not all lines, in $\mathbb{P}^{n}({\mathbb{C}})$.
Theorem 2.
Let $F_{i}$, $11\leq i\leq n+$ 1, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$. Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $11\leq i\leq n+$ 1, intersect transversally. Let $\Delta=(1-{1\over m_{1}})D_{1}+\cdots+(1-{1\over m_{n+1}})D_{n+1}$ be an orbifold divisor of $\mathbb{P}^{n}({\mathbb{C}})$. Assume that $\deg\Delta>n+1$. Then, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $W$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{1},\dots,m_{n+1}\}\geq\ell$ must be contained in $W$.
Note that $\mathbb{P}^{n}({\mathbb{C}})$ is a toric variety when we identify ${\mathbb{G}}_{m}^{n}$ with ${\mathbb{P}}^{n}({\mathbb{C}})\backslash\cup_{i=0}^{n}H_{i}$, where $H_{0},\dots,H_{n}$ are the $n+1$ coordinate hyperplanes. A toric variety $X$ is defined as a variety over ${\mathbb{C}}$ that contains ${\mathbb{G}}_{m}^{n}$ as a dense open subvariety, with the action of ${\mathbb{G}}_{m}^{n}$ on itself extends to an algebraic action of ${\mathbb{G}}_{m}^{n}$ on $X$. We refer to [5] for more details on toric varieties. Using the arguments in [13] and [6], we can extend our Main Theorem to projective toric varieties. Recall from [13] that an “admissible pair” is a couple $(X,V)$ where $V$ is a nonsingular variety embedded in a nonsingular projective variety $X$ in such a way that $D_{0}=X\setminus V$ is a normal crossings divisor. A Campana orbifold $(X,\Delta)$ is said to be associated to the admissible pair $(X,V)$ if $D_{0}$ is the support of $\Delta$.
Theorem 3.
Let $X$ be a nonsingular projective toric variety and let $D$ be an effective reduced divisor on the admissible pair $(X,\mathbb{G}_{m}^{n})$, such that $D$ is in general position with the boundary $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $A$ be a big divisor on $X$. For every $\epsilon>0$ there is a positive integral number $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
  1. (i)
    $N_{f}(D,r)-N_{f}^{(1)}(D,r)\leq_{\rm exc}\epsilon T_{A,f}(r)$, and
  2. (ii)
    $T_{D,f}(r)\leq_{\rm exc}N_{f}^{(1)}(D,r)+\epsilon T_{A,f}(r)$
for any non-constant entire orbifold curve $f$ on $(X,\Delta)$ such that the image of $f$ is not contained in $Z$.
Theorem 3 implies Campana’s conjecture for an orbifold pair that is a finite covering of a toric variety, as follows.
Theorem 4.
Let $X$ be a nonsingular projective toric variety, and let $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $Y$ be a nonsingular complex projective variety with a finite morphism $\pi:Y\to X$. Let $H:=\pi^{*}(D_{0})$ and $R\subset Y$ the ramification divisor of $\pi$ omitting components from the support of $H$. Assume that $\pi(R)$ and $D_{0}$ are in general position on $X$. Then there exists a positive integer $\ell$ such that if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
In Section 2, we prove a version of the Main Theorem for rational maps and explain the motivation for using parabolic Riemann surfaces. Section 3 reviews basic results from Nevanlinna theory relevant to parabolic settings, as well as the key theorems from [2]. Furthermore, we extend the GCD theorem developed in [14] and in [9] to parabolic setting. In Section 4, we establish the abc theorem for parabolic Riemann surfaces, building on the methods developed in [7]. Section 5 presents the proof of the Main Theorem, while Section 6 extends results to toric varieties.
2. Motivation through the rational case
In [7], a version of Vojta’s abc conjecture for algebraic tori over algebraic function fields is established. As we will see, the algebraic reduction method presented in this article is a key technical component in proving the Main Theorem. The core idea behind the proof is also inspired by the function field approach. In this section, we will prove a version of the Main Theorem for rational functions and explain the role of parabolic Riemann surfaces in this context.
2.1. The Rational Case
The statements and arguments apply verbatim to algebraic function fields. Here, let $K=\mathbb{C}(t)$ be the function field of $\mathbb{P}^{1}({\mathbb{C}})$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. We denote by ${\mathcal{O}}_{S}^{*}$ the set of $S$-units, which consists of rational functions with no zeros and poles outside of $S$. For a rational function $f=\frac{P}{Q}$, where $P$ and $Q$ are coprime polynomials, we define $\deg f:=\max\{\deg P,\deg Q\}$ and denote by ${\rm ord}_{\mathbf{p}}^{+}(f)$ the order of the zero of $f$ at $\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})$. We use the notation $N_{S}(f)$ (respectively $N_{S}^{(1)}(f)$) to count the number of zeros of $f$ with multiplicity (respectively, without multiplicity) outside of $S$. When $S$ is empty, we simply write $N(f)$ (respectively $N^{(1)}(f)$) to count the number of zeros with multiplicity (respectively, without multiplicity) on $\mathbb{P}^{1}(\mathbb{C})$.
Recall a simplified version of the main technical theorem in [7].
Theorem 5 ([7, Theorem 4]).
Let $G\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be a non-constant homogeneous polynomial with neither monomial factors nor repeated factors. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}$ such that, for all $\mathbf{u}:=(u_{1},\ldots,u_{n})\in({\mathcal{O}}_{S}^{*})^{n}\setminus Z$, we have either
  1. (i)
    the inequality
    $\max_{1\leq i\leq n}\{\deg(u_{i})\}\leq c_{0}\max\{1,|S|-2\}$
    holds,
  2. (ii)
    or the following two statements hold
    1. (a)
      $N_{S}(G(\mathbf{u}))-N_{S}^{(1)}(G(\mathbf{u}))\leq\epsilon\max_{1\leq i\leq n}\{\deg u_{i}\}$ if $G(u_{1},\ldots,u_{n})\neq 0$.
    2. (b)
      If $G(0,\ldots,0)\neq 0$ and $\deg_{X_{i}}G=\deg G=d$ for $1\leq i\leq n$, then
      $N_{S}^{(1)}(G(\mathbf{u}))\geq(\deg G-\epsilon)\max_{1\leq i\leq n}\{\deg u_{i}\}.$
Here $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
We will demonstrate how to apply the above theorem to show the following version of Main Theorem for rational case.
Theorem 6.
Let $G$ be a non-constant homogeneous polynomial in $\mathbb{C}[x_{0},\cdots,x_{n}]$ with no monomial factors and no repeated factors. Assume $[G=0]$ is in general position with the coordinate hyperplanes $H_{i}=[x_{i}=0]$, $0\leq i\leq n$. Then, for any $\epsilon>0$, there exist effectively computable positive integers $\ell$ and a proper Zariski closed subset $Z$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the following two inequalities hold:
  • (i)
    $N(G(\mathbf{f}))-N^{(1)}(G(\mathbf{f}))\leq\epsilon\deg\mathbf{f}$, and
  • (ii)
    $N^{(1)}(G(\mathbf{f}))\geq(\deg G-\epsilon)\deg\mathbf{f}$,
for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$, and for all $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$. Here, $\deg\mathbf{f}=\max\{\deg f_{0},\ldots,\deg f_{n}\}$.
In addition to Theorem 5, we also require the following result to prove Theorem 6.
Proposition 7 ([7, Corollary 6]).
Let $G=\sum_{\mathbf{i}\in I_{G}}a_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be a non-constant polynomial, where $a_{\mathbf{i}}\neq 0$ if $\mathbf{i}\in I_{G}$. Assume that $G(0,\dots,0)\neq 0$. Let $Z$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in I}a_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$ where $I$ is a non-empty subset of $I_{G}$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. Then, for all $(u_{1},\ldots,u_{n})\in(\mathcal{O}_{S}^{*})^{n}\setminus Z$, we have
$\sum_{\mathbf{p}\in S}{\rm ord}_{\mathbf{p}}^{+}(G(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|S|-2\}$
where $\tilde{c}=\frac{1}{2}\binom{n+\deg G}{n}(\binom{n+\deg G}{n}+1)$.
Proof of Theorem 6.
Denote by $\ell$ a sufficiently large integer to be determined later. Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros such that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$. Let
(2.1) $\tilde{S}_{\mathbf{f}}:=\{\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})\,|\,{\rm ord}_{\mathbf{p}}(f_{i})>0\quad\text{for some $0\leq i\leq n$}\}\cup\{\infty\}.$
Then
(2.2) $|\tilde{S}_{\mathbf{f}}|\leq\sum_{i=0}^{n}N_{S}^{(1)}(f_{i})+1\leq\frac{n+1}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}+1,$
and $f_{i}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$ for $0\leq i\leq n$. Let $\epsilon>0$ be given. We now apply Theorem 5 to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ independent of $\tilde{S}_{\mathbf{f}}$ of $\mathbb{A}^{n}$ such that we have either the inequality
(2.3) $\max_{0\leq i\leq n}\{\deg f_{i}\}(=\deg\mathbf{f})\leq c_{0}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\}$
holds or the following two inequalities hold:
(2.4) $N_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
(2.5) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$, where $\tilde{Z}_{0}$ is the projective closure of $Z_{0}$ in $\mathbb{P}^{n}({\mathbb{C}})$. Here $G_{0}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$. Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $Z_{0}$ do not depend on $\mathbf{f}$ as well. We first show that (2.3) does not hold if $\ell>2c_{0}n$. Indeed, if (2.3) holds with $\ell>2c_{0}n$, then it follows from (2.2) and the choice of $\ell$ that (2.3) yields
$\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\frac{(n+1)c_{0}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}<\frac{1}{2}\max_{0\leq i\leq n}\{\deg f_{i}\}+c_{0}.$
Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since not all the polynomials $f_{i}$ are constant, the assumption that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$ implies that $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
We continue to let $\ell>2c_{0}n$. As we have shown that (2.3) does not hold, it follows that (2.4) and (2.5) both hold if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$. Since $N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))$, it follows trivially from (2.5) that
(2.6) $N^{(1)}(G(\mathbf{f}))\geq N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
This shows (ii) holds.
It remains to show (i). We now compute zeros for points in $\tilde{S}_{\mathbf{f}}$. We will use Propostion 7. Let $W$ be the Zariski closed subset of $\mathbb{P}^{n}(\mathbb{C})$ containing all hypersurfaces defined by all possible subsums (including $G$) in the expansion of $G$. Assume that $\mathbf{f}(\mathbb{C})\not\subset W$. Since $G_{0}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
where $\tilde{c}=\frac{1}{2}\binom{n+degG}{n}(\binom{n+\deg G}{n}+1)$. Together with (2.2) and that $\ell\leq\max_{0\leq i\leq n}\{\deg f_{i}\}$, it yields
$m_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$
provided $\ell>(n+1)\tilde{c}_{0}\epsilon^{-1}.$ Combining with (2.4), we have
(2.7) $N(G_{0}(u_{1},\ldots,u_{n}))-N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
We now repeat the above arguments for $(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})$, $0\leq i\leq n$. Then we find a Zariski closed subset $Z$ and a positive constant $c$, independent of $\mathbf{f}$, such that for each $0\leq i\leq n$ we have
(2.8) $N\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)-N^{(1)}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)\leq 2\epsilon\deg\mathbf{f}.$
Since $f_{0},\ldots,f_{n}$ are polynomials without common zeros, for each $z\in\mathbb{C}$ there exists at least one $f_{i}$ such that $f_{i}(z)\neq 0$. Then
${\rm ord}^{+}_{z}(G(f_{0},\ldots,f_{n}))={\rm ord}_{z}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right).$
This shows that
$N(G(\mathbf{f}))-N^{(1)}(G(\mathbf{f}))$ $\leq\sum_{i=0}^{n}\left(N\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)-N^{(1)}\left(G(\frac{f_{0}}{f_{i}},\cdots,\frac{f_{n}}{f_{i}})\right)\right)$
[…]
The assertion (i) can be derived by adjusting $\epsilon$. ∎
2.2. Motivation for the proof of the Main Theorem
As in Section 2.1, the starting point of the proof of Theorem 6 is as the follows: For a given rational map, we define a set $S$ (depending on the map) that contains the zeros and poles of each coordinate function of the map. In this way, all the coordinate functions become $S$-units, and so we can apply Theorem 5. Importantly, the constants and the exceptional set obtained in Theorem 5 do not depend on $S$, allowing this method to be used for all rational maps. To apply this view point to the analytic case, we consider $f=(f_{0},\ldots,f_{n})$, where $f_{i}$, $0\leq i\leq n$, are entire functions without common zeros. Let $\mathcal{E}_{f}$ denote the set of all the zeros of $f_{i}$, $0\leq i\leq n$, and define $\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f}$. Then ${\mathcal{Y}_{f}}$ is a parabolic open Riemann surface. On ${\mathcal{Y}_{f}}$, the restriction $\tilde{f}_{i}:=f_{i}|_{\mathcal{Y}_{f}}$ becomes a unit on ${\mathcal{Y}}$ (i.e. it has neither zeros nor poles). This perspective was also adopted in [2], where they laid important foundational work. For our purposes, we will need to develop an analogue of Theorem 5 over parabolic open Riemann surfaces.
3. Nevanlinna theory and the GCD theorem on parabolic Riemann surfaces
3.1. Parabolic Riemann surfaces
[…]
1. For a hypersurface $D=\{Q=0\}$ in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q$ is a homogeneous polynomial of degree $d$, the proximity function is defined as
[…]
where $\lambda_{D}$ is the Weil-function on ${\mathbb{P}}^{n}({\mathbb{C}})$ given by
$\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~{}~{}{\bf x}\in\mathbb{P}^{n}({\mathbb{C}}).$
2. The characteristic (order) function is defined by
$T_{f,{\mathcal{Y}}}(r)=\int_{1}^{r}{dt\over t}\int_{B(t)}f^{*}\omega_{FS}.$
3. The counting function and the $k$-truncated counting function are defined by
$N_{f,{\mathcal{Y}}}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}{\rm ord}_{z}f^{*}D{dt\over t},~{}~{}~{}~{}~{}~{}~{}\mbox{and}~{}~{}N_{f,{\mathcal{Y}}}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
4. The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},{\mathcal{Y}}}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [14].
Parabolic First Main Theorem. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$ and $f:{\mathcal{Y}}\rightarrow\mathbb{P}^{n}({\mathbb{C}})$ be a holomorphic map with $f({\mathcal{Y}})\not\subset D$. Then
$T_{f,{\mathcal{Y}}}(r)=m_{f,{\mathcal{Y}}}(D,r)+N_{f,{\mathcal{Y}}}(D,r)+O(1).$
It is important to note that, when we write $O(g(x))$ for some function $g$, we mean that $Cg(x)$ where $C$ is a constant independent of ${\mathcal{Y}}$ throughout the paper.
Fix a nowhere vanishing global holomorphic vector field $\xi\in\Gamma({\mathcal{Y}},T_{\mathcal{Y}})$ (we denote it by ${\partial\over\partial z}$) on ${\mathcal{Y}}$. Such a vector field exists because ${\mathcal{Y}}$ is non-compact and consequently its holomorphic tangent bundle $T_{\mathcal{Y}}$ is holomorphically trivial. From Lemma 4.4 in [10],
$\int_{S_{r}}\log\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}=-\int_{1}^{r}\chi_{\sigma}(t)\frac{dt}{t}+2\varsigma\log r+O(1),$
where $\chi_{\sigma}(r)$ is the Euler characteristic of $B(r)$, and
[…]
which is independent of $r$ for $r$ being sufficiently large. Denote by
$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S(r)}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
[…]
(3.1) $\hat{\tau}:=\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
Note that $\hat{\tau}$ takes values in $[0,+\infty)$, is continuous on ${\mathcal{Y}}$ and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$. By the Poincare-Lelong formula, it is clear that
$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{r_{j}\over 2}\sum_{j=1}^{\infty}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
is a distribution of order 00 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$. The smoothing process was done in [2] as follows: as in the Appendix in [2], let
$h(r)=\begin{cases}0&\text{ for $r\leq{3\over 4}$}\\ {1\over 4\pi}{1\over 1+e^{{r-1\over(r-1)^{2}-1/16}}}&\text{for ${3\over 4}<r<{5\over 4}$}\\ {1\over 4\pi}&\text{for $r\geq{5\over 4}$}\end{cases}$
and
$H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside $[{1\over 2},{3\over 2}]$. Note that $H$ is an absolute function independent of $a_{j}$. This function $H(r)$ is smooth of $\log^{+}r$ with $0\leq H(r)-\log^{+}r\leq\log^{+}{3\over 2}<{1\over 2}$ and $H(r)=\log^{+}r$ when $H(r)\geq\log{3\over 2}$. Define
$\tau:=\log\sigma:=H(z)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right).$
Then $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
$\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right),$
where $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}):=\{z\in{\mathcal{Y}}:{1\over 2}r_{j}\leq|z-a_{j}|\leq{3\over 2}r_{j}\}$ are pairwise disjoint annuli. For $z\not\in\cup_{j=1}^{\infty}\overline{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}$, $\log^{+}\left|{r_{j}\over z-a_{j}}\right|=H\left(\left|{r_{j}\over z-a_{j}}\right|\right)=0$ for each $j$. Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are eqaual when $\tau(z)\geq\log{3\over 2}$. Let $B_{r}^{\sigma}:=\{z\in{\mathcal{Y}}~{}|~{}\sigma(z)<r\}$ be the $\sigma$-ball of radius $r$. Then the above discussion implies that
(3.2) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~{}~{}~{}~{}(t\geq 1)~{}~{}~{}\mbox{and}~{}~{}~{}~{}~{}{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}(t\geq{3\over 2}).$
We now show that the smooth 2-form $dd^{c}\sigma$ defines an order 0 distribution of finite mass on ${\mathcal{Y}}$ (see Lemma 3.2 in [2]) so that Parabolic Logarithmic Derivative Lemma (Lemma 9) holds for $\sigma$. We use the formula that, in polar coordinates $z=a+re^{i\theta}$ for some $a\in{\mathbb{C}}$, for a smooth function $\phi$, one has
$d^{c}\phi={1\over 4\pi}\left(r{\partial\phi\over\partial r}d\theta-{1\over r}{\partial\phi\over\partial\theta}dr\right).$
The smooth 2-form $dd^{c}\sigma$ is supported on $U$ since $\log\sigma$ is harmonic elsewhere. On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinates $z=re^{i\theta}$ one has (see the Appendix in [2])
$dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=O(1)dr\wedge d\theta=O(1)rdr\wedge d\theta$
since ${1\over 2}\leq r\leq{3\over 2}$ is bounded. Thus $dd^{c}H(|z|)$ is of finite mass on $A(0,{1\over 2},{3\over 2})\backslash{\mathcal{E}}$. On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$,
$dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$ $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)$
$=d\left(-r_{j}h\left({r_{j}\over r}-c\right)d\theta\right)=O(r_{j})dr\wedge d\theta=O(1)rdr\wedge d\theta.$
Since $\sum_{j=1}^{\infty}r^{2}_{j}\leq\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure. We conclude that $dd^{c}\log\sigma$ is of finite mass.
Using the above computations with ${\mathcal{E}}:=f^{-1}(D)$, we recall the following important result due to Z. Chen, D.T. Huynh, R. Sun and S.Y. Xie [2].
Proposition 8 ([2]).
Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$. Let ${\mathcal{Y}}_{f}:={\mathbb{C}}\backslash f^{-1}(D)$ with the smooth exhaustion function $\tau:=\log\sigma$ given above. Then
  1. (i)
    $0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),$
    for $r\geq 1$, where $\widetilde{f}$ the restriction of $f$ on $\mathcal{Y}_{f}$.
  2. (ii)
    For $r>1$,
    $\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r,$
    where $C$ is a constant independent of $f$ and $D$.
We outline a proof here.
Proof.
Let ${\mathcal{E}}:=f^{-1}(D)$ and write :=${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$. Denote by $B_{r}^{\sigma}$ and $S_{r}^{\sigma}$ the $\sigma$-ball of radius $r$ and its boundary. From (3.2), $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ ($t\geq 1$) and
${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}~{}~{}~{}~{}(t\geq{3\over 2}).$
Thus
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
$\leq\int_{1}^{r}{dt\over t}\int_{\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}+O(1)$
$=\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~{}~{}~{}~{}(r\geq 1).$
Recall in (3.1) that the radius $r_{j}>0$ can be chosen arbitrarily small. For our purpose, for each $j\geq 1$, we choose $r_{j}>0$ sufficiently small so that $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}$. Hence the above estimate yields
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),~{}~{}~{}~{}(r\geq 1).$
This proves (i).
To prove (ii), using the above computations and following the argument in [16, Proposition 3.3, pp. 32-33] (see also [2, Line 3 on page 12]), the weighted Euler characteristic satisfies
$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)=\int_{1}^{r}\#\{j:|a_{j}|<t\}{dt\over t}+O(\log r).$
Note that #$\#\{{\mathbb{D}}_{t}\cap{\mathcal{E}}\}$ is exactly $n_{f}^{(1)}(D,t)$, which is finite. Since $B^{\sigma}_{r}\subset{\mathbb{D}}_{r}$ for $r\geq 1$,
$\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r.$
This completes the proof of (ii). ∎
3.2. Nevanlinna theory on parabolic Riemann surfaces
Let ${\mathcal{Y}}$ be a parabolic open Riemann surface. The following is the starting point of the parabolic Nevanlinna theory.
Lemma 9 (Parabolic Logarithmic Derivative Lemma, Theorem 3.7 in [16] and Proposition 5.1 in [10]).
Let $f$ be a meromorphic function on ${\mathcal{Y}}$. For any $\delta>0$, one has
$m_{f^{\prime}/f,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{f,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
In particular, it gives the following
Lemma 10.
If $u$ is a nowhere vanishing holomorphic function on $\mathcal{Y}$ ((((i.e. $u$ is a unit)))), then, for any $\delta>0$, one has
$T_{u^{\prime}/u,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
We recall the following theorem which is an extension of H. Cartan’s theorem. It was from Theorem 6.3 in [10], although the statement is slightly different.
Theorem 11 (Theorem 6.3 in [10]).
Let $f=(f_{0},\dots,f_{n}):{\mathcal{Y}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspaces of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are holomorphic functions on ${\mathcal{Y}}$ without common zeros. Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$. Then, for $\epsilon>0$,
$\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}}(f)d\mu_{r}+N_{W,\mathcal{Y}}(0,r)$
$\leq_{\rm exc}(n+1+\epsilon)T_{f,{\mathcal{Y}}}(r)+\left(\frac{n(n+1)}{2}+2\epsilon\right)(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+O(1),$
where $W$ is the Wronskian of $f$, and the maximum is taken over all subsets $J$ such that $H_{j},j\in J$, are in general position.
The following theorem can be derived from the proof of [18, Theorem 2.1] and Lemma 9.
Theorem 12.
Let $f_{0},\dots,f_{n}$ be entire functions on $\mathcal{Y}$ without common zeros. Assume that $f_{n+1}$ is the holomorphic function on ${\mathcal{Y}}$ such that $f_{0}+\dots+f_{n}+f_{n+1}=0$. If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then, for any $\epsilon>0$, we have
$(1-\epsilon)T_{f_{j}/f_{i},{\mathcal{Y}}}(r)\ \leq_{\operatorname{exc}}\sum_{i=0}^{n+1}N_{f_{i},{\mathcal{Y}}}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for any pair $0\leq i,j\leq n$.
3.3. The moving GCD theorem for parabolic Riemann surfaces
One of the key components in [9] is the extension of the GCD theorem in [14] to the case where the coefficients of two coprime polynomials are allowed to be meromorphic functions of the same growth. In this subsection, we aim to extend this line of argument to the setting of parabolic Riemann surfaces. Before we proceed, let us briefly outline the development of the GCD theorem.
The foundational arguments in [14] are based on results initially established in [13] for the number field case. In [22], the GCD results concerning $\mathbb{G}_{m}^{n}$ from [13] were generalized to more general varieties by combining the approach in [21] with Diophantine approximation results for arbitrary varieties introduced by the first named author and Vojta [20]. More recently, a further generalization was achieved in [11], which applied Vojta’s results on integral points on semiabelian varieties for the number field case, as well as [15] for the complex case.
To proceed, we will first derive a version of the second main theorem with moving targets of the same growth, which will serve our purpose.
Let $\mathcal{Y}$ be a parabolic open Riemann surface. Let $a_{0},\ldots,a_{n}$ be meromorphic functions, not all zeros, on $\mathcal{Y}$, and let $L:=a_{0}X_{0}+\dots+a_{n}X_{n}$. Then $L$ defines a hyperplane $H$ in $\mathbb{P}^{n}$, over the field generated by $a_{0},\ldots,a_{n}$ over $\mathbb{C}$. We note that $H(z)$ is the hyperplane determined by the linear form $L(z)=a_{0}(z)X_{0}+\dots+a_{n}(z)X_{n}$ for $z\in\mathcal{Y}$ where $a_{0},\cdots,a_{n}$ are entire and there is no common zero of $a_{0},\cdots,a_{n}$ at $z$. The definition of the Weil function, proximity function and counting function can be easily extended to moving hyperplanes. For example,
(3.3) $\lambda_{H(z)}(P)=-\log\frac{|a_{0}(z)x_{0}+\cdots+a_{n}(z)x_{n}|}{\max\{|x_{0}|,\dots,|x_{n}|\}\max\{|a_{0}(z)|,\dots,|a_{n}(z)|\}},$
for $z\in\mathcal{Y}$ which is not a common zero of $a_{0},\cdots,a_{n}$, or a pole of any $a_{k}$, $0\leq k\leq n$. The First Main Theorem for a moving hyperplane $H$ can be stated as
(3.4) $T_{\mathbf{f},\mathcal{Y}}(r)=N_{\mathbf{f},\mathcal{Y}}(H,r)+m_{\mathbf{f},\mathcal{Y}}(H,r)+T_{\mathbf{a},\mathcal{Y}}(r)+O(1),$
where ${\mathbf{a}}:=[a_{0}:\cdots:a_{n}]$.
Without loss of generality, we will normalize the linear forms $L_{j}$, $1\leq j\leq q$, such that for each $1\leq j\leq q$, there exists $0\leq j^{\prime}\leq n$ with $a_{ij^{\prime}}=1$. Let $t$ be a positive integer and let $V(t)$ be the complex vector space spanned by the elements
$\left\{\prod a_{jk}^{n_{jk}}:n_{jk}\geq 0,\,\sum n_{jk}=t\right\},$
where the products and sums run over $1\leq j\leq q$ and $0\leq k\leq n$. Let $1=b_{1},b_{2},\cdots,b_{u}$ be a basis of $V(t)$ and $b_{1},\cdots,b_{w}$ a basis of $V(t+1)$. It’s clear that $u\leq w$. Moreover, we have
(3.5) $\liminf_{t\to\infty}\dim V(t+1)/\dim V(t)=1.$
Definition 13.
Let $V$ be a $\mathbb{C}$-vector space spanned by finitely many meromorphic functions. We say that a holomorphic map ${\mathbf{f}}=(f_{0},\dots,f_{n}):\mathcal{Y}\to\mathbb{P}^{n}({\mathbb{C}})$ is linearly non-degenerate over $V$ if whenever we have a linear combination $\sum_{i=1}^{m}a_{i}f_{i}=0$ with $a_{i}\in V$, then $a_{i}=0$ for each $i$. Otherwise we say that ${\mathbf{f}}$ is linearly degenerate over $V$.
The following formulation of the second main theorem with moving targets follows from the strategy in [23] and the proof of [17, Theorem A6.2.4], by adding the Wronskian term and computing the error term explicitly when applying the second main theorem (Theorem 11).
Theorem 14.
[…]
(3.6) $\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}(z)}({\mathbf{f}}(z))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.7) $\leq_{\operatorname{exc}}\left(\frac{w}{u}(n+1)+\epsilon\right)T_{\mathbf{f},\mathcal{Y}}(r)+\frac{w}{u}(n+1)(t+2)\max_{1\leq j\leq q}T_{L_{j},{\mathcal{Y}}}(r)$
(3.8) $+{1\over 2}(n+1)^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where $T_{L_{j},{\mathcal{Y}}}(r)=T_{[a_{j0}:\cdots:a_{jn}],{\mathcal{Y}}}(r)$, $w=\dim V(t+1)$, $u=\dim V(t)$, and the maximum is taken over all subsets $J$ of $\{1,\dots,q\}$ such that $H_{j}(z)$, $j\in J$, are in general position.
Using the above theorem, and through the same argument from the proofs of [14, Theorem 5.7], [8, Theorem 26], and [9, Theorem 4.4], we can get the following GCD theorem (note we now consider the case that $u_{1},\ldots,u_{n}$ are units on $\mathcal{Y}$). We first fix some notations: Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. For every positive integer $t$, we denote by $V_{F,G}(t)$ the (finite-dimensional) $\mathbb{C}$-vector space spanned by $\prod_{\alpha}\alpha^{n_{\alpha}}$, where $\alpha$ runs through all non-zero coefficients of $F$ and $G$, $n_{\alpha}\geq 0$ and $\sum n_{\alpha}=t$; we also put $d_{t}:=\dim V_{F,G}(t)$. For every integer $m\geq d$, we let $M=M_{m,n,d}=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$. We now define the characteristic function of a polynomial over $K$. Let $Q=\sum_{{\mathbf{i}}\in I_{Q}}a_{\mathbf{i}}\cdot{\mathbf{x}}^{\mathbf{i}}\in K[x_{1},\ldots,x_{n}]$, where ${\mathbf{i}}=(i_{1},\ldots,i_{n})$, ${\mathbf{x}}^{\mathbf{i}}=x_{1}^{i_{1}}\ldots x_{n}^{i_{n}}$, and $a_{\mathbf{i}}\neq 0$ if ${\mathbf{i}}\in I_{Q}$. We let
(3.9) $\|Q\|_{z}:=\max_{{\mathbf{i}}\in I_{Q}}\{|a_{\mathbf{i}}(z)|\},$
and define the characteristic function of $Q$ as
(3.10) $T_{Q,\mathcal{Y}}(r):=T_{[\cdots:a_{\mathbf{i}}:\cdots],\mathcal{Y}}(r),$
where $a_{\mathbf{i}}$ is taken for every ${\mathbf{i}}\in I_{Q}$.
Theorem 15 (GCD theorem on parabolic Riemann surfaces evaluating at units).
Let $n\geq 2$ be a positive integer. Let $u_{1},\ldots,u_{n}$ be units on $\mathcal{Y}$, i.e. holomorphic functions on $\mathcal{Y}$ without zeros and let ${\bf u}=(u_{1},\ldots,u_{n})$. Let $F,G$ be coprime polynomials in $n$ variables of the same degree $d>0$ over $K$. Assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Then there exists a positive constant $c$ and a positive integer $b$ such that if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, then, for any $\delta>0$, we have
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},{\mathcal{Y}}}(r)$
$+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where $w=d_{Mb}$, $u=d_{M(b-1)}$, $c^{\prime}=\frac{w}{u}M(b+2)+1$, and $c^{\prime\prime}=M\frac{w^{2}}{u}$, and $M^{\prime}=\binom{m+n}{n}-M$.
We outline a proof here.
Proof.
As in the proof of [9, Theorem 4.3], we assume that the ideal $(F,G)$ is proper in $K[x]$. Let $(F,G)_{m}:=K[x]_{m}\cap(F,G)$, where we choose $\{\phi_{1},\dots,\phi_{M}\}$ to be a basis of the $K-$vector space $(F,G)_{m}$ consisting of elements of the form $F{\bf x}^{\bf i},G{\bf x}^{\bf i}$. Let $\Phi:=(\phi_{1},\dots,\phi_{M})$. For each $z\in{\mathcal{Y}}$, we can construct a basis $B_{z}$ of $V_{m}:=K[{\bf x}]_{m}/(F,G)_{m}$ with monomial representatives ${\bf x}^{{\bf i}_{1}},\dots,{\bf x}^{{\bf i}_{M^{\prime}}}$ as in the proof of Theorem 5.7 in [14]. Let $I_{z}:=\{{\bf i}_{1},\dots,{\bf i}_{M^{\prime}}\}$. For each ${\bf i}\not\in I_{z}$ with $|{\bf i}|=m$, we have
${\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\in(F,G)_{m}$
for some choice $c_{z,{\bf j}}\in K$. Then, for each such ${\bf i}$, there is a linear form over $K$ such that
$\tilde{L}_{z,{\bf i}}(\phi_{1},\dots,\phi_{M})={\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}.$
We note that since there are correspondingly only finitely many choices of a monomial basis of $V_{m}$, there are only finitely many choices of $c_{z,{\bf j}}$, even as $z$ ranges over all of $\mathbb{C}$. Following the explicit construction in the proof of [9, Theorem 4.3], we can find a nonzero element $c_{z}\in V_{F,G}(M)$ such that the normalized linear form $L_{z,{\bf i}}:=c_{z}\tilde{L}_{z,{\bf i}}$ satisfies
$L_{z,{\bf i}}(\phi_{1},\dots,\phi_{M})=c_{z}\left({\bf x}^{\bf i}+\sum_{{\bf j}\in I_{z}}c_{z,{\bf j}}{\bf x}^{\bf j}\right)$
and
(3.11) $T_{L_{z,{\bf i}},{\mathcal{Y}}}(r)\leq M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)),$
by [9, (4.12)]. Moreover, the explicit construction also demonstrates that there are only a finitely many choice for $c_{z}$, even though $z$ runs through all of $\mathbb{C}$
Following the steps in the proof of Theorem 4.3 in [9], we then obtain
(3.12) $\log|L_{z,{\bf i}}(\Phi({\bf u}))|-\log\|L_{z,{\bf i}}\|_{z}\leq\log|{\bf u}(z)^{\bf i}|+\log\|F\|_{z}+\log\|G\|_{z}+O(1).$
Consequently, we have
(3.13) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}(\log|L_{z,{\bf i}}(\Phi({\bf u})|-\log\|L_{z,{\bf i}}\|)d\mu_{r}\leq M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)+M(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+O(1).$
Also, from (3.3) in the proof of [14, Theorem 5.7], we have
(3.14) $\int_{S(r)}\log\|\Phi({\bf u})\|d\mu_{r}=T_{\Phi({\bf u}),{\mathcal{Y}}}(r)+N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+O(1).$
Since the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$, we have that $\Phi({\bf u}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{M-1}$ is linearly nondegenerate over $V_{F,G}(Mb)=V(b)$. We note that, when computing Weil functions, we may assume that the coefficients of $L_{z,{\bf i}}$ are entire functions without common zeros. By this construction, for each $z\in{\mathcal{Y}}$, the hyperplanes defined by evaluating the linear forms $L_{z,{\bf i}}$ at $z$ with $|{\bf i}|=m$ and ${\bf i}\not\in I_{z}$ are in general position. Additionally, we observe that (3.11) and (3.13) remain valid under this choice of linear forms. Applying Theorem 14 to $\Phi({\bf u})$ and the set of the linear forms $\{L_{z,{\bf i}}:|{\bf i}|=m,{\bf i}\not\in I_{z}\}$, we have the following inequality:
(3.15) $\int_{S(r)}\sum_{|{\bf i}|=m,{\bf i}\not\in I_{z}}\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))d\mu_{r}+\frac{1}{u}N_{W}(0,r)$
(3.16) $\leq_{\operatorname{exc}}\left(\frac{w}{u}M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+\frac{w}{u}M^{2}(b+2)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))$
(3.17) $+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
where
$\lambda_{L_{z,{\bf i}}}(\Phi({\bf u}))=-\log|L_{z,{\bf i}}(\Phi({\bf u})|+\log\|\Phi({\bf u})\|+\log\|L_{z,{\bf i}}\|+O(1).$
By combining (3.15) with (3.13), (3.14), we get
$MN_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)+\frac{1}{u}N_{W}(0,r)$
$\leq_{\operatorname{exc}}\left(\frac{w}{u}M-M+\epsilon\right)T_{\Phi({\bf u}),\mathcal{Y}}(r)+M^{\prime}mnT_{{\bf u},\mathcal{Y}}(r)$
$+\left(\frac{w}{u}M^{2}(b+2)+M\right)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+M^{2}{w^{2}\over u}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1),$
which proves the theorem. ∎
Remark.
Let $\epsilon$ be a sufficiently small positive number. By (3.5), we find $t$ such that $\frac{w}{u}\leq 1+\epsilon$. An explicit upper bound for $t$ can be found in the proof of [3, Proposition 5.3] in term of $\epsilon$ and the number of generator for $V(1)$, which is less than $qn$. More precisely, if $\epsilon<\frac{1}{16}$, then we can find
(3.18) $t\leq\frac{3qn}{\epsilon}\log\frac{1}{\epsilon}$
such that $\frac{w}{u}\leq 1+\epsilon$. Furthermore, for such $t$, we have
(3.19) $u:=\dim V(t)\leq\left(\frac{1}{\epsilon}\right)^{qn}.$
Consequently,
(3.20) $w:=\dim V(t+1)\leq(1+\epsilon)\left(\frac{1}{\epsilon}\right)^{qn}.$
Therefore, the constants appeared in (3.6) depend only on $\epsilon$ and the number of hyperplanes.
Theorem 16 (The key theorem).
Let $n\geq 2$ be a positive integer. Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F$ and $G$ be coprime polynomials in $K[x_{1},\ldots,x_{n}]$. Then, for any $\epsilon>0$, there exist a positive integer $m$ and positive reals $c_{i}$, $1\leq i\leq 4$, all depending only on $\epsilon$, $n$, and the degree of $F$ and $G$ such that for all $n$-tuple $(u_{1},\ldots,u_{n})$ of units on $\mathcal{Y}$ satisfying
(3.21) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq c_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have either
(3.22) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
(3.23) $N_{\rm gcd,\mathcal{Y}}(F(u_{1},\ldots,u_{n}),G(u_{1},\ldots,u_{n}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
Furthermore, if $\epsilon>0$ is sufficiently small, then we may choose $m=C_{1}\epsilon^{-1}$, where $C_{1}$ is a positive real independent of $\epsilon$.
Proof.
We note that $c_{i},1\leq i\leq 4,$ will be determined during the course of the proof. As in the proof of [9, Theorem 4.1], we will assume that one of the coefficients in each expansion of $F$ and $G$ is 1. Recall that $M:=M_{m}:=2\binom{m+n-d}{n}-\binom{m+n-2d}{n}$ and $M^{\prime}:=M^{\prime}_{m}:=\binom{m+n}{n}-M=O(m^{n-2})$.
Let $\epsilon>0$. To establish (3.22) or (3.23), we can assume that $\epsilon$ is sufficiently small. Based on the above estimates, we can choose a real $C_{1}\geq 1$ independent of $\epsilon$ and the $u_{i},1\leq i\leq n,$ such that $m=C_{1}\epsilon^{-1}\geq 2d$, and
(3.24) $\frac{M^{\prime}mn}{M}\leq\frac{\epsilon}{4}.$
By (3.5) we may then choose a sufficiently large integer $b\in\mathbb{N}$ such that
(3.25) $\frac{w}{u}-1\leq\frac{\epsilon}{4mn},$
where $w:=\dim_{{\bf k}}V_{F,G}(Mb)$ and $u:=\dim_{{\bf k}}V_{F,G}(Mb-M)$.
Let ${\bf u}=(u_{1},\ldots,u_{n})$. Suppose that the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly independent over $V_{F,G}(Mb+1)$. Then, by Theorem 15, we have
(3.26) $N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)$ $\leq_{\rm exc}\left(\frac{M^{\prime}}{M}+\frac{w}{u}-1+\frac{\epsilon}{M}\right)mn\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)$
(3.27) $+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1).$
Let $c_{1}=4c^{\prime\prime}\epsilon^{-1}$. Then (3.21) reads
$c^{\prime\prime}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(1)\leq\frac{\epsilon}{3}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r).$
Together with (3.24) and (3.25), we derive from (3.26) and the above inequality that
$N_{\rm gcd,{\mathcal{Y}}}(F({\bf u}),G({\bf u}),r)\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c^{\prime}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
Finally, if the set $\{u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}:i_{1}+\cdots+i_{n}\leq m\}$ is linearly dependent over $V_{F,G}(Mb+1)$, i.e. there is a non-trivial relation
(3.28) $\sum_{\mathbf{i}}\alpha_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}=0,$
where the sum runs over those $\mathbf{i}\in\mathbb{Z}_{\geq 0}^{n}$ and $|\mathbf{i}|\leq m$, and $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$ for each $\mathbf{i}$ such that $\alpha_{\mathbf{i}_{0}}\neq 0$ for some $\mathbf{i}_{0}$. Then we have
(3.29) $\sum_{\mathbf{i}\neq\mathbf{i}_{0}}\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}}=-1.$
Since $\alpha_{\mathbf{i}}\in V_{F,G}(Mb+1)$, we have
(3.30) $T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}},\mathcal{Y}}(r)\leq T_{\alpha_{\mathbf{i}},\mathcal{Y}}(r)+T_{\alpha_{\mathbf{i_{0}}},\mathcal{Y}}(r)\leq 2(Mb+1)(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
By Theorem 12 and (3.30), the relation (3.29) implies
$T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)\leq_{\operatorname{exc}}C_{2}\cdot(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $C_{2}=4\binom{n+m}{n}(Mb+1)$ and $c_{3}={\binom{n+m}{n}}\left({\binom{n+m}{n}}+1\right)$. Then
(3.31) $T_{\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)$ $\leq T_{\frac{\alpha_{\mathbf{i}}}{\alpha_{\mathbf{i}_{0}}}\mathbf{u}^{\mathbf{i}-\mathbf{i}_{0}},\mathcal{Y}}(r)+T_{\frac{\alpha_{\mathbf{i}_{0}}}{\alpha_{\mathbf{i}}},\mathcal{Y}}(r)$
(3.32) $\leq_{\rm exc}c_{2}(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c_{2}=C_{2}+2(Mb+1)$. ∎
For the convenience of later application, we state the following result for $n=1$, which can be easily obtained from the theory of resultant (see [19, Proposition 10]).
Proposition 17.
Let $K$ be a subfield of the field of meromorphic functions on $\mathcal{Y}$. Let $F,G\in K[x_{0},x_{1}]$ be nonconstant coprime homogeneous polynomials. Then there exists a positive constant $c$ such that for any holomorphic functions $g_{0}$ and $g_{1}$ on ${\mathcal{Y}}$ with no common zeros, the following holds:
$N_{\rm gcd,\mathcal{Y}}(F(g_{0},g_{1}),G(g_{0},g_{1}),r)\leq c(T_{F,{\mathcal{Y}}}(r)+T_{G,{\mathcal{Y}}}(r)).$
4. The abc theorem for parabolic Riemann surfaces
We use the above key theorem (Theorem 16) in the following manner: Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are holomorphic functions on ${\mathcal{Y}}$ without zeros (i.e. $u_{i}$ are the units on ${\mathcal{Y}}$). Let $\mathbb{K}_{\mathbf{u}}$ be the subfield of the meromorphic function field such that, for any element $a\in\mathbb{K}_{\mathbf{u}}$, there exists a nonnegative real constant $c_{a}$ (which may depend on $a$) satisfying
$T_{a,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{a}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}).$
Notably, in our application, the constant $c_{a}$ can be effectively computed. We will use Theorem 16 for polynomials in the ring $\mathbb{K}_{\mathbf{u}}[x_{1},\ldots,x_{n}]$. Furthermore, we observe that $u_{i}^{\prime}/u_{i}\in\mathbb{K}_{\mathbf{u}}$ for $1\leq i\leq n$, since
(4.1) $T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+2\log T_{u_{i},{\mathcal{Y}}}(r)$
by Lemma 10.
Let $\mathbf{x}:=(x_{1},\ldots,x_{n})$. For $\mathbf{i}=(i_{1},\ldots,i_{n})\in\mathbb{Z}^{n}$, we let $\mathbf{x^{i}}:=x_{1}^{i_{1}}\cdots x_{n}^{i_{n}}$ and $\mathbf{u^{i}}:=u_{1}^{i_{1}}\cdots u_{n}^{i_{n}}$. For a non-constant polynomial $F(\mathbf{x})=\sum_{\mathbf{i}}a_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]:=\mathbb{K}_{\mathbf{u}}[x_{1},\dots,x_{n}]$, we define
(4.2) $D_{\mathbf{u}}(F)(\mathbf{x}):=\sum_{\mathbf{i}}\frac{(a_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{\mathbf{u}^{\mathbf{i}}}\mathbf{x}^{\mathbf{i}}=\sum_{\mathbf{i}}\left(a_{\mathbf{i}}\cdot\sum_{j=1}^{n}i_{j}\frac{u_{j}^{\prime}}{u_{j}}\right)\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}].$
A direct computation shows that
(4.3) $F(\mathbf{u})^{\prime}=D_{\mathbf{u}}(F)(\mathbf{u}),$
and that the following product rule
(4.4) $D_{\mathbf{u}}(FG)=D_{\mathbf{u}}(F)G+FD_{\mathbf{u}}(G)$
holds for $F,G\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$.
The following lemma shares the same proof of Lemma 5.1 in [9].
Lemma 18.
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are the units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then $F:=D_{\mathbf{u}}(G)$ and $G$ are coprime in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ unless there exists a non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $\sum_{i=1}^{n}|m_{i}|\leq 2\deg G$ such that $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},{\mathcal{Y}}}\leq T_{G,{\mathcal{Y}}}(r).$
Proof.
The first step of the proof of [9, Lemma 5.1] shows that we may assume that $G$ is irreducible. Write $G=\sum_{\mathbf{i}}b_{\mathbf{i}}\mathbf{x}^{\mathbf{i}}\in\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$, which contains at least two distinct terms, since $G$ has no monomial factors. Then $D_{\mathbf{u}}(G)/G\in\mathbb{C}^{*}$ since $D_{\mathbf{u}}(G)$ is not zero and $\deg D_{\mathbf{u}}(G)=\deg G$. Comparing the coefficients of $G$ and $D_{\mathbf{u}}(G)$, for nonzero $b_{\mathbf{i}}$ and $b_{\mathbf{j}}$ in $\mathbb{K}_{\mathbf{u}}$ ($\mathbf{i}\neq\mathbf{j}$), we have from (4.2) that
$\frac{(b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}})^{\prime}}{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}=\frac{(b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}})^{\prime}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}},$
where $\mathbf{i}=(i_{1},\ldots,i_{n})$, and $\mathbf{j}=(j_{1},\ldots,j_{n})$. It implies that $\frac{b_{\mathbf{i}}\mathbf{u}^{\mathbf{i}}}{b_{\mathbf{j}}\mathbf{u}^{\mathbf{j}}}\in\mathbb{C}$. Therefore,
$T_{u_{1}^{i_{1}-j_{1}}\cdots u_{n}^{i_{n}-j_{n}},\mathcal{Y}}\leq T_{G,{\mathcal{Y}}}(r).$
As an application of the key theorem (Theorem 16) and using the above lemma, we obtain the following theorem.
Theorem 19.
Let $\mathbf{u}=(u_{1},\ldots,u_{n})$, where $u_{i}$, $1\leq i\leq n$, are units on ${\mathcal{Y}}$. Let $G$ be a non-constant polynomial in $\mathbb{K}_{\mathbf{u}}[\mathbf{x}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exists a positive integer $m$ and a positive constants $\tilde{c}_{1}$ and $\tilde{c}_{2}$ depending only on $\epsilon$, $n$, and the degree and the coefficients of $G$, but independent of $\mathbf{u}$, such that the following holds: If
(4.5) $\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)\geq\tilde{c}_{1}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
then we have either
(4.6) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2m$, or
$N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)\leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
Proof.
We first note that the constant ${\tilde{c}}_{1},{\tilde{c}}_{2}$ will be determined during the course of the proof. Let $z_{0}\in\mathcal{Y}$. If $v_{z_{0}}(G(\mathbf{u}))\geq 2$, then it follows from (4.3) that $v_{z_{0}}(D_{\mathbf{u}}(G)(\mathbf{u}))=v_{z_{0}}(G(\mathbf{u}))-1$. Hence,
$\mathop{\mathrm{min}}\{v_{z_{0}}^{+}(G(\mathbf{u})),v_{z_{0}}^{+}(D_{\mathbf{u}}(G)(\mathbf{u}))\}\geq v_{z_{0}}^{+}(G(\mathbf{u}))-\mathop{\mathrm{min}}\{1,v_{z_{0}}^{+}(G(\mathbf{u}))\}.$
Consequently,
(4.7) $N_{\gcd,{\mathcal{Y}}}(G(\mathbf{u}),D_{\mathbf{u}}(G)(\mathbf{u}),r)\geq N_{G(\mathbf{u}),{\mathcal{Y}}}(0,r)-N^{(1)}_{G(\mathbf{u}),{\mathcal{Y}}}(0,r).$
Note that there exists a nonnegative real $c_{G}$ such that
(4.8) $T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
since the coefficients of $G$ are in $\mathbb{K}_{\mathbf{u}}$. Furthermore,
(4.9) $T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)$ $\leq 2T_{G,{\mathcal{Y}}}(r)+\deg G\sum_{i=1}^{n}T_{u_{i}^{\prime}/u_{i},{\mathcal{Y}}}(r)$
(4.10) $\leq_{\rm exc}(2c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\}),$
and
(4.11) $T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)\leq_{\rm exc}(3c_{G}+n\deg G)(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
by (4.8) and (4.9). By Lemma 18 and (4.8), $G$ and $D_{\mathbf{u}}(G)$ are either coprime or
$T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}T_{G,{\mathcal{Y}}}(r)\leq_{\rm exc}c_{G}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\max_{1\leq i\leq n}\{\log T_{u_{i},{\mathcal{Y}}}(r)\})$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum_{i=0}^{n}|m_{i}|\leq 2\deg G$, i.e., either $G$ and $D_{\mathbf{u}}(G)$ are coprime or (4.6) holds.
It remains to consider where $G$ and $D_{\mathbf{u}}(G)$ are coprime. Let $\epsilon>0$ be given, which may assume to be sufficiently small. We apply Theorem 16 and find $m$ and $c_{i}>0$, $1\leq i\leq 4$, such that if (4.5) holds with $\tilde{c}_{1}\geq c_{1}$ , then we have either
(4.12) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}c_{2}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r))+c_{3}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for some non-trivial tuple of integers $(m_{1},\ldots,m_{n})$ with $|m_{1}|+\cdots+|m_{n}|\leq 2m$, or
(4.13) $N_{\rm gcd,\mathcal{Y}}(G({\mathbf{u}}),D_{\mathbf{u}}(G)({\mathbf{u}}),r)\leq_{\rm exc}\frac{\epsilon}{2}\max_{1\leq i\leq n}T_{u_{i},\mathcal{Y}}(r)+c_{4}(T_{G,{\mathcal{Y}}}(r)+T_{D_{\mathbf{u}}(G),{\mathcal{Y}}}(r)).$
We can conclude the proof by combining (4.11), (4.12), and (4.13), and letting
(4.14) $\tilde{c}_{1}=\max\{c_{1},2c_{4}(3c_{G}+n\deg G)\epsilon^{-1}\},~{}~{}~{}\mbox{and}~{}~{}\tilde{c}_{2}=\max\{c_{G},c_{3}+3c_{2}c_{G}+nc_{2}\deg G\}.$
The proof of [7, Theorem 4] can be adapted to the current situation, leading to the following result.
Theorem 20.
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $Z$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all $n$-tuple $\mathbf{u}=(u_{1},\ldots,u_{n})$ of units on ${\mathcal{Y}}$ with $\mathbf{u}(\mathcal{Y})\not\subset Z$ and
(4.15) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have
(4.16) $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$
if $G(\mathbf{u})\neq 0$. Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$, but not $\mathcal{Y}$ and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
We outline a proof here.
Proof.
We follow the proof of [7, Theorem 4]. Let $u_{i},1\leq i\leq n$, be units on ${\mathcal{Y}}$. We apply Theorem 19 through several reduction steps to derive a contradiction. Assume that (4.16) does not hold and let $c_{0}\geq\tilde{c}_{1}$, where $\tilde{c}_{1}$ is the constant from (4.5) in Theorem 19. Then, there exists a positive integer $M_{1}$, depending only on $\epsilon$, $n$, and the degree of $G$, but independent of $\mathbf{u}$, such that
(4.17) $T_{u_{1}^{m_{1}}\cdots u_{n}^{m_{n}},\mathcal{Y}}(r)\leq_{\rm exc}\tilde{c}_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O\left(\max_{1\leq j\leq n}\{\log T_{u_{j},{\mathcal{Y}}}(r)\}\right)$
for a non-trivial $n$-tuple $(m_{1},\ldots,m_{n})$ of integers with $\sum|m_{i}|\leq M_{1}$. Thus, we have
(4.18) $\lambda_{1}:={u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}}\in\mathbb{K}_{\mathbf{u}}.$
We may assume $\gcd(m_{1},\ldots,m_{n})=1$. By Lemma 14 in [7], $(m_{1},\ldots,m_{n})$ extends to a basis $(m_{1},\ldots,m_{n})$, $(a_{21},\ldots,a_{2n}),\ldots,(a_{n1},\ldots,a_{nn})$ of $\mathbb{Z}^{n}$ with
(4.19) $|a_{i1}|+\cdots+|a_{in}|\leq M_{1}+n\quad\text{ for }2\leq i\leq n.$
Consider the change of variables
(4.20) $\Lambda_{1}:=x_{1}^{m_{1}}\cdots x_{n}^{m_{n}},\quad\text{and }\quad X_{1,i}:=x_{1}^{a_{i1}}\cdots x_{n}^{a_{in}}\quad\text{for }2\leq i\leq n$
and put
(4.21) $\beta_{1,i}=u_{1}^{a_{i1}}\cdots u_{n}^{a_{in}}\quad\text{for }2\leq i\leq n.$
Let $A_{1}$ denote the $n\times n$ matrix whose rows are the above basis of $\mathbb{Z}^{n}$. Then we formally express the above identities as
(4.22) $(\Lambda_{1},X_{1,2},\ldots,X_{1,n})=(x_{1},\ldots,x_{n})^{A_{1}}\quad\text{and}\quad(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})=(u_{1},\ldots,u_{n})^{A_{1}}.$
Let $B_{1}=A_{1}^{-1}$. The entries of $B_{1}$ can be bounded from above in terms of $M_{1}$ and $n$. We have
(4.23) $(x_{1},\ldots,x_{n})=(\Lambda_{1},X_{1,2},\ldots,X_{1,n})^{B_{1}}\quad\text{and}\quad(u_{1},\ldots,u_{n})=(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})^{B_{1}}.$
Let $G_{1}(\Lambda_{1},X_{1,2},\ldots,X_{1,n})\in{\mathbb{C}}[\Lambda_{1},X_{1,2},\ldots,X_{1,n}]$ with no monomial factors and
(4.24) $G((\Lambda_{1},X_{1,2},\ldots,X_{1,n})^{B_{1}})=\Lambda_{1}^{d_{1}}X_{1,2}^{d_{2}}\cdots X_{1,n}^{d_{n}}G_{1}(\Lambda_{1},X_{1,2},\ldots,X_{1,n})$
for some integers $d_{i}$, $1\leq i\leq n$. Since the transformations in (4.22) and (4.23) are inverses of each other, and $G$ has no repeated irreducible factors, it follows that $G_{1}$ also has no repeated irreducible factors. The coefficients of $G_{1}$ are the same as the coefficients of $G$ and $\deg(G_{1})$ can be bounded from above explicitly in terms of $M_{1}$, $n$, and $\deg(G)$. Consider $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})\in\mathbb{C}(\lambda_{1})[X_{1,2},\ldots,X_{1,n}]\subset{\mathbb{K}_{\mathbf{u}}}[X_{1,2},\ldots,X_{1,n}]$.
For the particular change of variables in (4.22), (4.23), and (4.24) (that depends on the matrix $A_{1}$), we apply the Lemma 16 in [7] to find a nonconstant polynomial $H_{1}^{\prime}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $G_{1}(\lambda_{1},X_{1,2},\ldots,X_{1,n})$ has neither monomial nor repeated irreducible factors if $H_{1}^{\prime}(u_{1},\ldots,u_{n})\not\equiv 0$. It is important to note that the polynomial $H_{1}^{\prime}$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$. We now take $H_{1}$ to be the product of all such $H_{1}^{\prime}$ where $A_{1}$ ranges over the finitely many elements of $\operatorname{GL}_{n}(\mathbb{Z})$ with $\|A_{1}\|_{\infty}\leq M_{1}+n$. From Lemma 16 in [7]  we know that $\deg H_{1}$ depends only on $\epsilon$, $n$ and $\deg G$.
Since the $u_{i}$’s, $\lambda_{1}$, and $\beta_{1,j}$’s are the units on $\mathcal{Y}$, we have
(4.25) $\begin{split}N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N_{G(\mathbf{u}),\mathcal{Y}}^{(1)}(0,r)=N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n}),\mathcal{Y}}^{(1)}(0,r).\end{split}$
On the other hand, by taking $\tilde{c}_{1}\geq 2\tilde{c}_{2}\epsilon^{-1}$, we deduce from (4.15) and (4.17) that
(4.26) $T_{{u_{1}}^{m_{1}}\cdots{u_{n}}^{m_{n}},\mathcal{Y}}(r)\ \leq_{\rm exc}\epsilon\max_{1\leq j\leq n}\{T_{u_{j},{\mathcal{Y}}}(r)\}.$
Then, by (4.22), (4.23) and (4.26), we have:
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}={\rm O}\left(\max\{T_{\lambda_{1},\mathcal{Y}}(r),T_{\beta_{1,2},\mathcal{Y}}(r),\ldots,T_{\beta_{1,n},\mathcal{Y}}(r)\}\right)={\rm O}\left(\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\right).$
More precisely, we find positive real numbers $b_{1,1}$ and $b_{1,2}$ such that
(4.27) $b_{1,1}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\},\quad\text{and }\quad\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\leq b_{1,2}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
In conclusion, at the end of this step we have that
$\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}\geq b_{1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds, and the inequality
(4.28) $N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}(0,r)-N_{G_{1}(\lambda_{1},\beta_{1,2},\ldots,\beta_{1,n})}^{(1)}(0,r)$ $\leq_{\rm exc}\frac{\epsilon}{b_{1,2}}\max_{2\leq i\leq n}\{T_{\beta_{1,i},\mathcal{Y}}(r)\}$
fails to hold under the assumption that $H_{1}(u_{1},\ldots,u_{n})\not\equiv 0$.
The above is the first step. There are $(n-1)$ many steps in total as in the proof of [7, Theorem 4]. Let $12\leq s\leq n-$ 1 and suppose that we have completed the step $s-1$. This includes the construction of $H_{s-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$ with degree depends on $\epsilon$, $n$ and $\deg G$ only. The last one is Step $(n-1)$ resulting in $H_{n-1}\in{\mathbb{C}}[x_{1},\ldots,x_{n}]$. We now define $H=H_{1}\cdots H_{n-1}$. Then $\deg H$ depends only on $\epsilon$, $n$ and $\deg G$, since each $H_{i}$ does so. Moreover, the polynomial $H$ is not only independent of $\mathbf{u}$, but also independent of $\mathcal{Y}$.
Suppose $H(u_{1},\ldots,u_{n})\not\equiv 0$. Assume we go through all the $(n-1)$ steps to get the polynomial
$P(X_{n-1,n}):=G_{n-1}(\lambda_{1},\ldots,\lambda_{n-1},X_{n-1,n})\in{\mathbb{K}_{\mathbf{u}}}[X_{n-1,n}]$
such that its degree can be bounded explicitly in terms of $M_{n-1},\ldots,M_{1}$, $n$, and $\deg(G)$. At the end of the step $(n-1)$, we have that $\beta:=\beta_{n-1,n}$, which is a (non-constant) unit on ${\mathcal{Y}}$, such that
(4.29) $T_{\beta,\mathcal{Y}}(r)\geq b_{n-1,1}c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
holds and the inequality
(4.30) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{\epsilon}{b_{n-1,2}}T_{\beta,\mathcal{Y}}(r)$
fails to hold. Here, $b_{n-1,1}$ and $b_{n-1,2}$ are computable positive real numbers. On the other hand, since $H_{n-1}(u_{1},\ldots,u_{n})\not\equiv 0$, the polynomial $P(X_{n-1,n})$ has neither monomial nor repeated irreducible factors, and thus, similar to (4.7), we have
(4.31) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\geq N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r).$
By Proposition 17, there is a positive constant $C_{1}$ independent of $\beta$ such that
(4.32) $N_{\gcd,{\mathcal{Y}}}(P(\beta),D_{\beta}(P)(\beta),r)\leq C_{1}(T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)).$
Similar to the arguments for (4.9), we find a positive constant $C_{2}$ independent of $\beta$ such that
(4.33) $T_{P,{\mathcal{Y}}}(r)+T_{D_{\beta}(P),{\mathcal{Y}}}(r)\leq_{\rm exc}C_{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+O(\log T_{\beta,\mathcal{Y}}(r)).$
Then by (4.29), (4.31), (4.32), and (4.33), we have
(4.34) $N_{P(\beta)}(0,r)-N^{(1)}_{P(\beta)}(0,r)\leq_{\rm exc}\frac{C_{1}C_{2}}{b_{n-1,1}c_{0}}T_{\beta,\mathcal{Y}}(r)+O(\log T_{\beta,\mathcal{Y}}(r)),$
which contradicts (4.30) by taking $c_{0}\geq 2C_{1}C_{2}b_{n-1,2}b_{n-1,1}^{-1}\epsilon^{-1}$. ∎
Theorem 21.
Let $G=\sum_{\mathbf{i}\in I_{G}}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}\in$ be a non-constant polynomial in $\mathbb{C}[x_{1},\cdots,x_{n}]$. Assume that $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$. Let $Z_{1}$ be the Zariski closed subset that is the union of hypersurfaces of $\mathbb{A}^{n}$ of the form $\sum_{\mathbf{i}\in J}\alpha_{\mathbf{i}}{\mathbf{x}}^{\mathbf{i}}=0$, where $J$ is a non-empty subset of $I_{G}$. Let $u_{1},\ldots,u_{n}$ be the units on $\mathcal{Y}$ such that the image of ${\bf u}:=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ is not contained in $Z_{1}$. Then
(4.35) $N_{G({\bf u}),\mathcal{Y}}(0,r)\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$.
Proof.
Since $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$, we may assume without loss of generality that $G(0,\ldots,0)=1$, and write
(4.36) $G(\mathbf{u})=1+\sum_{1\leq i\leq n}\alpha_{\mathbf{i}_{i}}u_{i}^{d}+\sum_{\mathbf{i}\in I_{G}\setminus I}\alpha_{\mathbf{i}}{\mathbf{u}}^{\mathbf{i}}$
where $\alpha_{\mathbf{i}_{i}}\neq 0$ for $0\leq i\leq n$ and $I=\{\mathbf{i}_{0}:=(0,0,\ldots,0),\mathbf{i}_{1}:=(d,0,\ldots,0),\ldots,\mathbf{i}_{n}:=(0,\ldots,0,d)\}$. If ${\mathbf{u}}({\mathbb{C}})$ is not contained in $Z_{1}$, we may use Theorem 12 to show that
$(d-\epsilon)\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\leq N_{G(\mathbf{u}),\mathcal{Y}}(0,r)+c(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
where $c=\frac{1}{2}\binom{n+d}{n}(\binom{n+d}{n}+1)$. ∎
Theorem 22 (The abc Theorem on parabolic Riemann surfaces).
Let $G$ be a non-constant polynomial in $\mathbb{C}[x_{1},\ldots,x_{n}]$ with no monomial factors and no repeated factors. Then, for any $\epsilon>0$, there exist a positive real number $c_{0}$ and a proper Zariski closed subset $W$ of $\mathbb{A}^{n}(\mathbb{C})$ such that for all holomorphic maps $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$ and satisfying
(4.37) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}\geq c_{0}(\mathfrak{X}^{+}_{\sigma}(r)+\log r),$
we have the following two statements hold
  1. (a)
    $N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq_{\rm exc}\epsilon\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}$.
  2. (b)
    If $G(0,\ldots,0)\neq 0$ and $\deg_{x_{i}}G=\deg G=d$ for $1\leq i\leq n$, then
    $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\rm exc}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
Here, $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $W$ can be expressed as the zero locus of a finite set $\Sigma\subset K[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ (not on $\mathcal{Y}$) and can be determined explicitly, (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
Proof.
The proof follows from Theorem 20 and Theorem 6.4. Let $W=Z\cup Z_{1}\cup[G=0]$, where $Z$ and $Z_{1}$ are the exceptional set from Theorem 20 and Theorem 6.4 for $\frac{\epsilon}{2}$ respectively. Let $\mathbf{u}=(u_{1},\ldots,u_{n}):\mathcal{Y}\to(\mathbb{C}^{*})^{n}$ with $\mathbf{u}(\mathcal{Y})\not\subset W$. Theorem 20 implies that
$N_{G(\mathbf{u}),\mathcal{Y}}(0,r)-N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\leq\frac{\epsilon}{2}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
Together with (4.35) in Theorem 6.4, we derive
(4.38) $N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}-c(\mathfrak{X}^{+}_{\sigma}(r)+\log r).$
By replacing the constant $c_{0}$ in (4.37) by $c_{1}=\max\{c_{0},2c\epsilon^{-1}\}$, we may derive from (4.38) that
$N^{(1)}_{G(\mathbf{u}),\mathcal{Y}}(0,r))\geq_{\operatorname{exc}}(\deg G-\epsilon)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}}(r)\}.$
5. Proof of the main theorem
Proof of the Main Theorem.
Let $f=(f_{0},\ldots,f_{n}):\mathbb{C}\mapsto\mathbb{P}^{n}(\mathbb{C})$, where $f_{0},\ldots,f_{n}$ are entire functions without common zero. We may assume $f_{i}$ is not identically zero by letting $Z$ contain all coordinate hyperplanes. When $f$ is rational, the assertion follows directly from Theorem 6, as in this case we have
$N_{G(f)}(0,r)=N(G(f))\log r,\quad N^{(1)}_{G(f)}(0,r)=N^{(1)}(G(f))\log r,\quad\text{ and }T_{f}(r)=(\deg f)\log r$
for $r$ sufficiently large. Therefore, it remains to consider the situation when $f$ is not rational. In this case, we have
$\log r=o(T_{f}(r)).$
Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes. We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$. By (ii) of Proposition 8,
(5.1) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r).$
Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$. Thus
(5.2) $\mathfrak{X}_{\sigma}^{+}(r)\leq\frac{n+1}{\ell}T_{f}(r)+O(\log r),$
where $\ell$ is a positive integer to be determined. Let $\tilde{f}_{i}$ be the restriction of $f_{i}$ to $\mathcal{Y}_{f}$ and let $u_{i}:=\frac{\tilde{f}_{i}}{\tilde{f}_{0}}$. Then $u_{i},1\leq i\leq n$, are units on $\mathcal{Y}_{f}$. Denote by ${\tilde{u}}=(1,u_{1},\ldots,u_{n})$ and $\mathbf{u}=(u_{1},\ldots,u_{n})$. Then $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}=T_{\tilde{u},\mathcal{Y}_{f}}(r)+O(1)=T_{{\tilde{f}},\mathcal{Y}_{f}}(r)+O(1),$ where $\tilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}$.
We verify that
(5.3) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)$
holds, where $c_{0}$ is a sufficient large constant determined in Theorem 22. Indeed, if it is not true, then there exists a constant $c>0$ and a sequence $r_{n}\rightarrow+\infty$ such that
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r_{n})\}\leq c(\mathfrak{X}_{\sigma}^{+}(r_{n})+\log r_{n}).$
Using (ii) of Proposition 8, (5.2) and the fact that $\log r=o(T_{f}(r))$, we have that
$T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})\leq{c(n+2)\over\ell}T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})+o(T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})),$
which gives a contradiction when $\ell>c(n+2)$. Hence (5.3) holds. Therefore, by Theorem 22, there exist a proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ and a positive constant $c$, which are independent of $f$, such that the following two statements hold
(5.4) $N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
(5.5) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
Here, $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$. Therefore, $\tilde{G}({\mathbf{u}})=G(\tilde{u})$. Since $\widetilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}=\mathbb{C}\setminus{f^{-1}(D)}$ and $\tilde{f}_{i}$ has no zero on $\mathcal{Y}_{f}$, it is obvious that
(5.6) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)=N^{(1)}_{G(\widetilde{f}),\mathcal{Y}_{f}}(0,r)\leq N^{(1)}_{G(f)}(0,r).$
Then it follows immediately from (5.5) that
$N^{(1)}_{G(f)}(0,r)\geq_{exc}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq(\deg G-\epsilon)T_{f}(r)$
by (ii) of Proposition 8 and the fact that $\log r=o(T_{f}(r))$. This proves (ii).
To show (i), we apply (5.6) again and obtain the following:
$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq N_{G(f)}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))$
$=N_{G(f)}(0,r)-N_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r)+N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r))$
$\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\},$
by the First Main Theorem, Theorem 6.4, (5.4) and (5.5). Again as $\log r=o(T_{f}(r))$, it yieids
(5.7) $N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)\leq\epsilon T_{f}(r)$
by (i) of Proposition 8. Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
6. Toric Varieties
Proof of Theorem 3.
We recall the following setup of finding a natural finite open covering of $X$ from the proof of [13, Theorem 4.4]. Let $\Sigma$ be the fan corresponding to the smooth projective toric variety $X$. Then there is a finite affine covering $\{X_{\eta}\}$ of $X$, where $\eta\in\Sigma$ is a n-dimensional smooth cone with an isomorphism $i_{\eta}:X_{\eta}\to\mathbb{A}^{n}$. This isomorphism restricts to an automorphism of $\mathbb{G}_{m}^{n}$, where we identify $\mathbb{G}_{m}^{n}\subset X_{\eta}$ naturally as a subset of $X$ and $\mathbb{G}_{m}^{n}\subset\mathbb{A}^{n}$ in the standard way such that $\mathbb{A}^{n}\setminus\mathbb{G}_{m}^{n}$ consists of the affine coordinate hyperplanes $\{x_{i}=0\}$, $1\leq i\leq n$. Under this identification, an entire orbifold curve $f$ into $(X,\Delta)$ corresponds to a collection of $n$-tuples of meromorphic functions ${i}_{\eta}(f)=(u_{\eta,1},\ldots,u_{\eta,n})$ such that the zero multiplicity of each $u_{\eta,i}$, $1\leq i\leq n$, is at least $\ell$ if not zero. If $z_{0}\in\mathbb{C}$ is a pole of some $u_{\eta,i}$, $1\leq i\leq n$, then $f(z_{0})$ is in a component of $D_{0}$, the boundary of $X$, with multiplicity
$-\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
Let $\widetilde{{i}_{\eta}}:X_{\eta}\to\mathbb{P}^{n}({\mathbb{C}})$ be the composition of $i_{\eta}$ with the natural embedding of $\mathbb{A}^{n}$ to $\mathbb{P}^{n}({\mathbb{C}})$. Then $\widetilde{{i}_{\eta}}(f)=(f_{\eta,0},\ldots,f_{\eta,n})=[1:u_{\eta,1}:\cdots:u_{\eta,n}]:\mathbb{C}\to\mathbb{P}^{n}({\mathbb{C}})$, where $f_{\eta,0},\ldots,f_{\eta,n}$ are entire functions with no common zero. Furthermore, the pullback $(\widetilde{{i}_{\eta}}^{-1})^{*}(D|_{X_{\eta}})$ of $D$ to $\mathbb{P}^{n}({\mathbb{C}})$ is defined by some nonconstant homogeneous polynomial $F_{\eta}\in\mathbb{C}[x_{0},\ldots,x_{n}]$ with no monomial and repeated factors and $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$ since $D$ is reduced and is in general position with the boundary of $\mathbb{G}_{m}^{n}$ in $X$. By [25, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
(6.1) $T_{\widetilde{{i}_{\eta}}(f)}(r)\leq c_{\eta,A}T_{A,f}(r)+O(1)$
if the image of ${i}_{\eta}(f)$ is not contained in $Z_{\eta,A}$. Let $M$ be the number of $X_{\eta}$ in $\Sigma$. We may now apply the Main Theorem to each $X_{\eta}$. Let $\epsilon>0$. Then there exists a proper Zariski closed subset $W_{\eta}\subset X_{\eta}$ for each $\eta\in\Sigma$ such that for sufficiently large integer $\ell$, we have
(6.2) $N_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)-N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\frac{\epsilon}{Mc_{\eta,A}}T_{\widetilde{{i}_{\eta}}(f)}(r),$
and
(6.3) $N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\geq_{\rm exc}(\deg F_{\eta}-\frac{\epsilon}{Mc_{\eta,A}})T_{\widetilde{{i}_{\eta}}(f)}(r),$
for all $f$ such that the image of $f$ restricting to $X_{\eta}$ is not contained in $W_{\eta}$. Since $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$, it implies that $\deg F_{\eta}\cdot T_{\widetilde{{i}_{\eta}}(f)}(r)=T_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(r)+O(1).$ Then by the first main theorem and (6.3), we have
(6.4) $m_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\frac{\epsilon}{Mc_{\eta,A}}T_{\widetilde{{i}_{\eta}}(f)}(r).$
Let $W$ be the union of $W_{\eta}$ and $Z_{\eta,A}$ for $\eta\in\Sigma$. Then
(6.5) $N_{f}(D,r)-N_{f}^{(1)}(D,r)\leq\sum_{\eta\in\Sigma}N_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)-N^{(1)}_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\epsilon T_{A,f}(r),$
and
(6.6) $m_{f}(D,r)\leq\sum_{\eta\in\Sigma}m_{F_{\eta}(\widetilde{{i}_{\eta}}(f))}(0,r)\leq_{\rm exc}\epsilon T_{A,f}(r)$
for all orbifold curve $f:\mathbb{C}\to(X,\Delta)$ with multiplicity at least $\ell$ along $D_{0}$ and $f(\mathbb{C})\not\subset W$. Finally, the assertion (ii) follows from (6.5) and the following inequality:
$N_{f}(D,r)=T_{D,f}(r)-m_{f}(D,r)\geq_{\rm exc}T_{D,f}(r)-\epsilon T_{A,f}(r)$
by (6.6). ∎
Proof of Theorem 4.
Let $\pi:Y\to X$ be a finite morphism. We recall from [4, Lemma 1] and the proof of [6, Theorem 6] that $R$ is a big divisor since $(Y,\Delta)$ is of general type. Let $R_{0}$ be an irreducible component of $R$. Without loss of generality, we let $R_{0}=R$. Otherwise, we simply repeat the following steps for each irreducible component. Let $D=\pi(R)$, which is in general position with $D_{0}$ by assumption. Now, consider $f:\mathbb{C}\to(Y,\Delta)$ as a nonconstant orbifold entire curve. Since $\pi^{*}D$ has multiplicity at least 2 along $R$, we have
(6.7) $N_{f}(R,r)\leq N_{\pi\circ f}(D,r)-N^{(1)}_{\pi\circ f}(D,r).$
We are now in a position to apply Theorem 3 for a small $\epsilon>0$ (to be determined later) and an ample divisor $A$ on $X$. Consequently, there exists a positive integer $\ell_{1}$ and a proper Zariski closed subset $W_{1}$ of $X$ such that if the orbifold $(X,\Delta_{0})$ has multiplicities at least $\ell_{1}$ along $D_{0}$, then
(6.8) $N_{\pi\circ f}(D,r)-N^{(1)}_{\pi\circ f}(D,r)\leq_{\operatorname{exc}}\epsilon T_{A,\pi\circ f}(r),$
and
(6.9) $T_{D,\pi\circ f}(r)\leq_{\rm exc}N_{\pi\circ f}^{(1)}(D,r)+\epsilon T_{A,\pi\circ f}(r)$
if the image of $\pi\circ f$ is not contained in $W_{1}$. Assuming that $\ell\geq\ell_{1}$ and the image of $f$ is not contained in $\pi^{-1}(W_{1})$, then by (6.7) and (6.8), we have
(6.10) $N_{f}(R,r)\leq_{\operatorname{exc}}\epsilon T_{A,\pi\circ f}(r).$
Since $R\leq\pi^{*}(D)$ (as divisors), the functorial property implies that
(6.11) $m_{f}(R,r)\leq m_{\pi\circ f}(D,r)+O(1)=T_{D,\pi\circ f}(r)-N_{\pi\circ f}(D,r)+O(1)$
Together with (6.9), we have:
(6.12) $m_{f}(R,r)\leq\epsilon T_{A,\pi\circ f}(r)+O(1).$
Combining the two equations (6.10) and (6.12) and using the functorial property of characteristic functions, we have:
(6.13) $T_{R,f}(r)\leq 2\epsilon T_{\pi^{*}A,f}(r)+O(1).$
Since $R$ is a big divisor on $Y$, by [25, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that
(6.14) $T_{\pi^{*}A,f}(r)\leq cT_{R,f}(r)+O(1),$
if the image of $\mathbf{f}$ is not contained in $W_{0}$. Combining (6.14) with (6.13), we obtain
$T_{R,f}(r)\leq_{\operatorname{exc}}2c\epsilon T_{R,f}(r)+O(1).$
This is not possible as $c$ is independent of $\epsilon$, and $\epsilon$ can be taken sufficiently small. In conclusion, we have found a sufficiently large integer $\ell$ such that if $(Y,\Delta)$ has multiplicities at least $\ell$ along $D_{0}$, then the image of any orbifold entire curve $f$ on $(Y,\Delta)$ is contained in $\pi^{-1}(W_{1})\cup W_{0}$. ∎